OpenStudy (mikezack123):

Which graph represents the solution set for the following compound inequality? 7 > −2x + 3 ≥ − 7

OpenStudy (mikezack123):

OOOPPSSSS

OpenStudy (mikezack123):

! SECOND PLEASE

OpenStudy (mikezack123):

there u go sorry bout that

OpenStudy (baseballer2014):

\[7>2x+3\ge-7 \]subtract three from both sides \[4>2x \ge-10\] divide each side by 2\[2>x \ge-5\] so now look at the graphs which graph has -5 colored and 2 left open?

OpenStudy (mikezack123):

A.

OpenStudy (mikezack123):

wait what...?

OpenStudy (baseballer2014):

hang on that was my mistake

OpenStudy (mikezack123):

what??? I'm really confused now

OpenStudy (baseballer2014):

\[4>-2x \ge-10\] divide by -2 \[-2<x \le5\]

OpenStudy (anonymous):

sorry thats nt true

OpenStudy (baseballer2014):

the signs don't flip?

OpenStudy (anonymous):

when u multply inequilty with negative sign ... the equility sign flips automatically keep it in mind

OpenStudy (anonymous):

well go on @baseballer2014 u r going well

OpenStudy (mikezack123):

so the answer would be B.

OpenStudy (anonymous):

hey its not B

OpenStudy (mikezack123):

omg...

OpenStudy (mikezack123):

im extremley confused

OpenStudy (anonymous):

@baseballer2014 is going gd .. he did the corect one

OpenStudy (baseballer2014):

Mike the answer is B based on my work

OpenStudy (anonymous):

sorry @MikeZack123 is B

OpenStudy (mikezack123):

but its B?

OpenStudy (baseballer2014):

yes it is B

OpenStudy (baseballer2014):

@inkyvoyd